OpenStudy (anonymous):

A scientist measures the angle x and the distance y between the Earth and the Sun. explain how the scientist can use only these two measurements to calculate the distance between the Sun and the shooting star.

OpenStudy (anonymous):

I just need someone to tell me if the scientist would use tan, sin, or cosine ratio

OpenStudy (anonymous):

@NoelGreco @ILovePuppiesLol

OpenStudy (anonymous):

I think they would use tangent?

OpenStudy (fortytherapper):

Since he would have to sides, I would just find the third angle (Since the third angle would be found using 180-90-x), then sine

OpenStudy (fortytherapper):

two*

OpenStudy (anonymous):

I don't have to find the angle. I need to say if the scientist would use tan, sin, or cosine

OpenStudy (fortytherapper):

I would use sine if I was him

OpenStudy (anonymous):

So it wouldn't be tan?

OpenStudy (anonymous):

the distance between the sun and the shooting star is equal to the value of side b whitch is the hypotenuse.

OpenStudy (fortytherapper):

Because he has the right angle and the x angle himself, so he would know that a triangle equals 180 degrees I would then put on my lab goggles and calculate \[180 - 90 - x\] to find that third angle. I would lastly use the law of sines to find that vertical distance \[\frac{ Shooting Star Distance }{ \sin(x) } = \frac{ Side Y }{ The \angle I just found}\]

OpenStudy (fortytherapper):

After that, Phythagoras for hypotenuse

OpenStudy (anonymous):

now u know use ur sohcahtoa to find side b.

OpenStudy (fortytherapper):

I didn't say you did, I'm just saying what the scientist would do

OpenStudy (anonymous):

what trig identities is adjecent/hypothenus?

OpenStudy (anonymous):

cosine

OpenStudy (anonymous):

there you go

OpenStudy (anonymous):

Can you explain why they would use cosine?

OpenStudy (anonymous):

surething, costheta is adjecent over hypothenus, you are giving the adjecent side but the hypothenus whisc is the distance you are looking for is unknown, therefore you need to look for an identities that has adjecent and hypothenus.

OpenStudy (anonymous):

Okay makes sense. Thanks :)

OpenStudy (anonymous):